OpenStudy (anonymous):

Need help:2^(x+5) = 9 How do you get 2 and 9 to have the same bases? or how else could you do this problem?

OpenStudy (agent47):

take natural log on both sides: ln(2^(x+5))=ln9

OpenStudy (agent47):

use the property: lna^b=b*lna, so you get: ln(2^(x+5))=(x+5)ln2=ln9

OpenStudy (agent47):

(x+5)=ln9/ln2 (x+5)=ln9/ln2

OpenStudy (agent47):

x=(ln9/ln2)-5

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

do you know a good website that goes over natural logs and e well?

OpenStudy (anonymous):

because I'm trying to do this next problem, and I don't know how to do it :(. This is the question: lnx+ln(x−6)=2